# Regularized Langevin Dynamics for Combinatorial Optimization

### v1

@@ Abstract
This work proposes a simple yet effective sampling framework for combinatorial optimization (CO). Our method builds on discrete Langevin dynamics (LD), an efficient gradient-guided generative algorithm. However, we observed that directly applying LD often leads to limited exploration. To overcome this limitation, we propose the Regularized Langevin Dynamics (RLD), which enforces an expected distance between the sampled and current solutions, effectively avoiding local minima. We develop two CO solvers on top of RLD, one based on simulated annealing (SA) and the other one based on neural network (NN). Empirical results on three classical CO problems demonstrate that both of our methods can achieve comparable or better performance against the previous state-of-the-art (SOTA) SA and NN-based solvers. In particular, our SA algorithm reduces the running time of the previous SOTA SA method by up to 80%, while achieving equal or superior performance. In summary, RLD offers a promising framework for enhancing both traditional heuristics and NN models to solve CO problems.
@@ 1 Introduction
Combinatorial Optimization (CO) problems are central challenges in computer science and operations research (Papadimitriou & Steiglitz, 1998), with diverse real-world applications such as supply chain management, logistics optimization (Chopra & Meindl, 2001), workforce scheduling (Ernst et al., 2004), financial portfolio management (Rubinstein, 2002; Lobo et al., 2007), compiler optimization (Trofin et al., 2021; Zheng et al., 2022), and bioinformatics (Gusfield, 1997). Despite their wide-ranging utility, CO problems are inherently difficult due to their non-convex nature and often NP-hard complexity, making them intractable in polynomial time by exact solvers. Traditional CO algorithms often rely on hand-crafted, domain-specific heuristics, which are costly and difficult to design, posing significant challenges in solving novel or complex CO problems.
Recent advancements in neural network (NN)-based learning (Bengio et al., 2020) and simulated annealing (SA) (Kirkpatrick et al., 1983) algorithms have redefined approaches to combinatorial optimization by minimizing dependence on manual heuristics:
  • Neural Network Models: NN-based methods leverage reinforcement learning (Khalil et al., 2017; Qiu et al., 2022), unsupervised learning (Karalias & Loukas, 2020a; Wang et al., 2022; Wang & Li, 2023; Sanokowski et al., 2024) or generative models (Kool et al., 2019; Zhang et al., 2023; Sun & Yang, 2023; Li et al., 2023, 2024) to learn optimization strategies directly from data. By automating the process, these models replace handcrafted heuristics with learned representations and decision-making processes, enabling tailored solutions refined through training rather than manual adjustment.
@@
Discrete Langevin dynamics (LD) (Zhang et al., 2022; Sun et al., 2022) and the corresponding diffusion models (Chen et al., 2023; Austin et al., 2021) have greatly advanced the recent development of both NN and SA solvers. The key idea of LD is to guide the iterative sampling via the gradient, for a more efficient searching/generation process. For instance, DIFUSCO (Sun & Yang, 2023) adopts continuous diffusion models from computer vision to address the discrete nature of CO problems, outperforming previous end-to-end neural models in both accuracy and computational efficiency. Additionally, DiffUCO (Sanokowski et al., 2024) generalizes DIFUSCO by eliminating the need for labeled training data, using unsupervised learning for CO problems. Meanwhile, advanced SA-based CO solvers have demonstrated performance on par with state-of-the-art NN-based approaches. Sun et al. (2023) underscore the advantages of LD-based SA method, including their simplicity, superior speed-quality trade-offs, and generalizability to new CO problems, as they require no training or problem-specific customization. However, existing discrete LD/diffusion methods are all adapted from the methods (Welling & Teh, 2011; Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Ho et al., 2020; Song & Ermon, 2020) in the continuous domain, this raises important questions: Is there any difference between CO and continuous optimization? Do these adapted methods sufficiently consider the nature of discrete data? Exploring these questions is the central focus of this paper.
@@
Our key observation is that the optimization process is more prone to local optima in a discrete domain than in a continuous one. That is, local optima in a continuous domain typically has a zero gradient (under the smoothness condition) but this is often not true in a discrete domain, where the gradients may be very large in magnitude but pointing to an infeasible region. Such a difference makes the escaping of local optima more difficult in a discrete domain than in a continuous one, with the common strategy of adding a random noise as in LD. We propose to address this issue by enforcing a constant norm of the expected distance between the sampled solution and the current solution during the searching process. In other words, we control the magnitude of the update in LD, encouraging the search to explore more promising areas. We name this sampling method Regularized Langevin Dynamics (RLD). We apply RLD on both SA and NN-based CO solvers, leading to Regularized Langevin Simulated Annealing (RLSA) and Regularized Langevin Neural Network (RLNN). Our empirical evaluation on three CO problems demonstrate thes significant improvement of RLSA and RLNN over both SA and NN baselines. Notably, RLSA only needs 20% running time to outperform the previous SOTA SA baselines. It shows a clear efficiency advantage with either less or more searching steps.
@@
To summarize, we propose a new variant of discrete Langevin dynamics for CO by regularizing the expected update magnitude on the current solution at each step. Our method is featured by its simplicity, effectiveness, and wide applicability to both SA and NN-based solvers, indicating its strong potential in addressing CO problems.
@@ 2 Preliminary
2.1 Combinatorial Optimization Problems
@@
where $\beta>0$ is the penalty coefficient that should be sufficiently large, such that the minima of Equation 2 corresponds to the feasible solutions in Equation 1. $H(\mathbf{x})$ is also generally named as the energy function, and its associated energy-based model (EBM) is defined as
@@
where $\tau>0$ is the temperature controlling the smoothness of the distribution, and $Z=\sum_{\mathbf{x}\in\{0,1\}^{N}}\exp(-H(\mathbf{x})/\tau)$ is the normalization factor, typically intractable. When $\tau$ is small, the probability mass of $p_{\tau}$ tends to concentrate around low-energy samples, making the task of solving Equation 1 equivalent to sampling from $p_{\tau}(\mathbf{x})$. Markov Chain Monte Carlo (MCMC) (Lecun et al., 2006) is the most widely used method for sampling from the EBM defined above. However, directly applying MCMC may lead to inefficiencies due to the non-smoothness introduced by the small $\tau$. To mitigate this issue, the simulated annealing (SA) technique is commonly employed to gradually decrease $\tau$ towards zero during the MCMC process.
@@ 2.2 Langevin Dynamics
$\mathbf{x}^{\prime}=\mathbf{x}+\frac{\alpha}{2}s(\mathbf{x})+\sqrt{\alpha}\zeta,\quad\zeta\in\mathcal{N}(0,\mathbf{I}_{N\times N}),$ (4)
@@
Recently, Zhang et al. (2022) have extended LD to discrete space by rewriting Equation 4 as
$q(\mathbf{x}^{\prime}|\mathbf{x})=\frac{\exp(-{\frac{1}{2\alpha}\|\mathbf{x}^{\prime}-\mathbf{x}-\frac{\alpha}{2}s(\mathbf{x})}\|_{2}^{2})}{Z(\mathbf{x})},$ (5)
where $Z(\mathbf{x})$ is the normalization factor. For discrete data, the above distribution could be factorized coordinatewisely, i.e., $q(\mathbf{x}^{\prime}|\mathbf{x})=\prod_{i=1}^{N}q(\mathbf{x}^{\prime}_{i}|\mathbf{x})$, into a set of categorical distributions:
$q(\mathbf{x}^{\prime}_{i}|\mathbf{x})\propto\exp({\frac{1}{2}s(\mathbf{x})_{i}(\mathbf{x}_{i}^{\prime}-\mathbf{x}_{i})-\frac{(\mathbf{x}_{i}^{\prime}-\mathbf{x}_{i})^{2}}{2\alpha})}.$ (6)
When $\mathbf{x}$ is a binary vector, we can obtain the flipping (changing the value of $\mathbf{x}_{i}$ from 00 to $1$, or $1$ to 00) probability $q(\mathbf{x}^{\prime}_{i}=1-\mathbf{x}_{i}|\mathbf{x})$ as
$\texttt{Sigmoid}(\frac{1}{2}s(\mathbf{x})_{i}(1-2\mathbf{x}_{i})-\frac{1}{2\alpha}).$ (7)
In particular, it can be shown that the discrete Langevin sampler is a first-order approximation to the locally-informed proposal (Zanella, 2017) in the following form.
$q(\mathbf{x}^{\prime}_{i}|\mathbf{x})\propto\exp({\frac{1}{2}p(\mathbf{x}^{\prime})-\frac{1}{2}p(\mathbf{x})-\frac{(\mathbf{x}_{i}^{\prime}-\mathbf{x}_{i})^{2}}{2\alpha})}.$ (8)
@@ 3.1 Regularized Langevin Dynamics
How to select the step size $\alpha$ is critical to the effectiveness of LD for CO. In this work, we propose to select $\alpha$ by regularizing the update with the L2222 distance111there is no difference between L1111 and L2222 distances on binary data, we leave the choice on other discrete data as future study. between the sampled and current solutions:
$\begin{split}q(\mathbf{x}^{\prime}_{i}|\mathbf{x})\propto&\exp({\frac{1}{2}s(\mathbf{x})_{i}(\mathbf{x}_{i}^{\prime}-\mathbf{x}_{i})-\frac{(\mathbf{x}_{i}^{\prime}-\mathbf{x}_{i})^{2}}{2\alpha})},\\ \text{s.t.}\qquad&\mathbb{E}_{q(\mathbf{x}^{\prime}|\mathbf{x})}[\|\mathbf{x}^{\prime}-\mathbf{x}\|_{2}]=d,\end{split}$ (9)
where $d$ is a positive integer. When $\mathbf{x}$ is binary, we could explicitly write out the expectation in Equation 9 with the flipping probability $q(\mathbf{x}^{\prime}_{i}=1-\mathbf{x}_{i}|\mathbf{x})$ :
$\sum_{i=1}^{N}\texttt{Sigmoid}(\frac{1}{2}s(\mathbf{x})_{i}(1-2\mathbf{x}_{i})-\frac{1}{2\alpha})=d.$ (10)
We find this simple regularization method very effective in mitigating the local optima issue in CO, in the sense that the regularization term enforces the change of a fixed magnitude to the solution, regardless of the gradient. We name our method as Regularized Langevin Dynamics (RLD), and we proceed to introduce its applications in both SA and NN-based CO solvers.

### v2

@@ Abstract
This work proposes a simple yet effective sampling framework for combinatorial optimization (CO). Our method builds on discrete Langevin dynamics (LD), an efficient gradient-guided generative paradigm. However, we observe that directly applying LD often leads to limited exploration. To overcome this limitation, we propose the Regularized Langevin Dynamics (RLD), which enforces an expected distance between the sampled and current solutions, effectively avoiding local minima. We develop two CO solvers on top of RLD, one based on simulated annealing (SA), and the other one based on neural network (NN). Empirical results on three classic CO problems demonstrate that both of our methods can achieve comparable or better performance against the previous state-of-the-art (SOTA) SA- and NN-based solvers. In particular, our SA algorithm reduces the runtime of the previous SOTA SA method by up to 80%, while achieving equal or superior performance. In summary, RLD offers a promising framework for enhancing both traditional heuristics and NN models to solve CO problems. Our code is available at https://github.com/Shengyu-Feng/RLD4CO.
@@ 1 Introduction
Combinatorial Optimization (CO) problems are central challenges in computer science and operations research (Papadimitriou & Steiglitz, 1998), with diverse real-world applications such as logistics optimization (Chopra & Meindl, 2001), workforce scheduling (Ernst et al., 2004), financial portfolio management (Rubinstein, 2002; Lobo et al., 2007), distributed computing (Zheng et al., 2022; Feng et al., 2025a), and bioinformatics (Gusfield, 1997). Despite their wide-ranging utility, CO problems are inherently difficult due to their non-convex nature and often NP-hard complexity, making them intractable in polynomial time by exact solvers. Traditional CO algorithms often rely on hand-crafted domain-specific heuristics, which are costly and difficult to design, posing significant challenges in solving novel or complex CO problems.
Recent advances in simulated annealing (SA) (Kirkpatrick et al., 1983) and neural network (NN)-based learning (Bengio et al., 2020; Cappart et al., 2023) algorithms have redefined approaches to combinatorial optimization by minimizing dependence on manual heuristics:
@@
  • Neural Network Models: NN-based methods leverage supervised learning (Kool et al., 2019; Zhang et al., 2023; Sun & Yang, 2023; Li et al., 2023, 2024), reinforcement learning (Khalil et al., 2017; Qiu et al., 2022; Feng & Yang, 2025) or unsupervised learning (Karalias & Loukas, 2020; Wang et al., 2022; Wang & Li, 2023; Sanokowski et al., 2024) to learn optimization strategies directly from data. By automating the process, these models replace hand-crafted heuristics with learned representations and decision-making processes, enabling tailored solutions refined through training rather than manual adjustment.
@@
Langevin dynamics (LD) (Welling & Teh, 2011) and related diffusion models (whose inference is built on LD) (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song & Ermon, 2019; Song et al., 2020) have greatly advanced the recent development of SA- and NN-based solvers. The key idea of LD is to guide the iterative sampling via the gradient for a more efficient searching/generation process. This gradient-informed approach has been adapted to the discrete domain—examples include GWG (Grathwohl et al., 2021), PAFS (Sun et al., 2022), and the discrete Langevin sampler (Zhang et al., 2022)—and yields SA solvers that attain state-of-the-art (SOTA) results on various CO benchmarks (Sun et al., 2023). Meanwhile, discrete diffusion models have also shown substantial gains. For example, DIFUSCO (Sun & Yang, 2023) adopts continuous diffusion models from computer vision to address the discrete nature of CO problems, outperforming previous end-to-end neural models in both the solution quality and computational efficiency. Additionally, DiffUCO (Sanokowski et al., 2024) generalizes DIFUSCO by eliminating the need for labeled training data, using unsupervised learning for CO problems. However, existing discrete LD/diffusion methods are mostly adapted from the methods in the continuous domain, this raises important questions: Is there any difference between continuous optimization and combinatorial optimization? Do these adapted methods sufficiently consider the nature of discrete data? Exploring these questions is the central focus of this paper.
@@
Our key observation is that the optimization process is more prone to local optima in a discrete domain than in a continuous one. That is, local optima in a continuous domain typically have a zero gradient (under the smoothness condition), but this is often not true in a discrete domain, where the gradient may be very large in magnitude, but pointing to an infeasible region. Such a difference makes the escaping of local optima more difficult in a discrete domain than in a continuous one, with the common strategy of adding random noise as in LD. We propose to address this issue by enforcing a constant norm of the expected Hamming distance between the sampled solution and the current solution during the searching process. In other words, we control the magnitude of the update in LD, encouraging the search to explore more promising areas. We name this sampling method Regularized Langevin Dynamics (RLD). We apply RLD on both SA- and NN-based CO solvers, leading to Regularized Langevin Simulated Annealing (RLSA) and Regularized Langevin Neural Network (RLNN). Our empirical evaluation on three CO problems demonstrate the significant improvement of RLSA and RLNN over both SA and NN baselines. Notably, RLSA only needs 20% runtime to outperform the previous SOTA SA baselines. And it shows a clear efficiency advantage with either less or more sampling steps. Compared with previous diffusion models, RLNN could be efficiently trained with a local optimization objective in an unsupervised manner, eliminating the need for labeled data or the estimation of a long-term return throughout the sampling process.
@@
To summarize, we propose a new variant of discrete Langevin dynamics for combinatorial optimization by regularizing the expected update magnitude on the current solution at each step. Our method is featured by its simplicity, effectiveness, and wide applicability to both SA- and NN-based solvers, indicating its strong potential in addressing CO problems.
2 Related Work
2.1 Simulated Annealing for CO
Simulated Annealing (SA) (Metropolis et al., 1953; Hastings, 1970; Neal, 1996; IBA, 2001) is a classic metaheuristic for combinatorial optimization (WANG et al., 2009; Bhattacharya et al., 2014; Tavakkoli-Moghaddam et al., 2007; Seçkiner & Kurt, 2007; Chen & Ke, 2004). However, traditional SA methods often require up to millions of sampling steps for convergence, which limits its applicability on large‐scale instances. The recent advances in discrete Markov Chain Monte Carlo (MCMC) offer the potential solutions by accelerating the convergence with locally-balanced proposals (Zanella, 2017). For example, GWG (Grathwohl et al., 2021) uses a first‐order approximation to estimate energy changes within the $1$‐Hamming‐ball neighborhood, steering each proposal toward high‐density regions. To mitigate local‐optima issues caused by small neighborhoods, PAS/PAFS (Sun et al., 2022) instead samples $d$ coordinates sequentially at each step. Their SA variant iSCO (Sun et al., 2023) achieves SOTA performance on various CO benchmarks. To address the inefficiency of sequential sampling in PAS/PAFS, Zhang et al. (2022) derive a discrete Langevin sampler from the continuous domain, assigning gradient‐informed change probabilities to all coordinates in parallel. However, direct transformation of the discrete Langevin sampler into an SA framework overlooks key domain differences, resulting in severe local‐optima issues. In this work, we build on these discrete MCMC developments by proposing a simple method that unifies parallel sampling with effective local‐optima escape.
2.2 Neural Solvers for CO
The neural network (NN) models have recently garnered vast attention in solving CO problems (Bengio et al., 2020; Cappart et al., 2023; Feng et al., 2025b). The NN-based solvers could be roughly categorized into three classes according to the training methods, namely the supervised learning-based (Li et al., 2018a; Gasse et al., 2019; Sun & Yang, 2023; Li et al., 2023, 2024), reinforcement learning-based (Khalil et al., 2017; Qiu et al., 2022; Feng & Yang, 2025), and unsupervised learning-based (Karalias & Loukas, 2020; Wang et al., 2022; Wang & Li, 2023; Zhang et al., 2023; Sanokowski et al., 2024). Although neural networks offer strong expressivity, the end-to-end neural solvers still struggle with the inherent non-convexity of CO problems. To address this, recent works (Sun & Yang, 2023; Li et al., 2023, 2024) have introduced discrete diffusion models, leveraging their success in capturing complex, multimodal distributions from image generation. However, these diffusion approaches depend on high-quality training data—often expensive to generate for large-scale CO instances, sometimes requiring hours per instance. DiffUCO (Sanokowski et al., 2024) alleviates this by using unsupervised learning guided by energy-based models, but its training still relies on reinforcement learning across the entire sampling trajectory, leading to inefficiency. In contrast, our RLNN method could be optimized only through a local objective, dramatically improving training efficiency while achieving comparable performance to SOTA NN-based solvers.
@@ 3 Preliminary
3.1 Combinatorial Optimization Problem
@@
where $\beta>0$ is the penalty coefficient that should be sufficiently large, such that the minimum of Equation 2 corresponds to the feasible solution in Equation 1. $H(\mathbf{x})$ is also generally named as the energy function, and its associated energy-based model (EBM) is defined as
@@
where $\tau>0$ is the temperature controlling the smoothness of $p_{\tau}(\mathbf{x})$, and $Z=\sum_{\mathbf{x}\in\{0,1\}^{N}}\exp(-H(\mathbf{x})/\tau)$ is the normalization factor, typically intractable. When $\tau$ is small, the probability mass of $p_{\tau}$ tends to concentrate around low-energy samples, making the task of solving Equation 1 equivalent to sampling from $p_{\tau}(\mathbf{x})$. Markov Chain Monte Carlo (MCMC) (Lecun et al., 2006) is the most widely used method for sampling from the EBM defined above. However, directly applying MCMC may lead to inefficiencies due to the non-smoothness introduced by the small $\tau$. To mitigate this issue, the simulated annealing (SA) technique is commonly employed to gradually decrease $\tau$ towards zero during the MCMC process.
@@ 3.2 Langevin Dynamics
$\mathbf{x}^{\prime}=\mathbf{x}+\frac{\alpha}{2}s(\mathbf{x})+\sqrt{\alpha}\zeta,\quad\zeta\sim\mathcal{N}(0,\mathbf{I}_{N\times N}),$ (4)
@@
Recently, Zhang et al. (2022) have extended LD to a discrete Langevin sampler by rewriting Equation 4 as
$\begin{split}q(\mathbf{x}^{\prime}|\mathbf{x})&=\frac{\exp\bigl{(}-\frac{1}{2\alpha}\|\mathbf{x}^{\prime}-\mathbf{x}-\frac{\alpha}{2}s(\mathbf{x})\|_{2}^{2}\bigr{)}}{Z(\mathbf{x})}\\ &=\frac{\exp\bigl{(}\frac{1}{2}s(\mathbf{x})^{T}(\mathbf{x}^{\prime}-\mathbf{x})-\frac{1}{2\alpha}\|\mathbf{x}^{\prime}-\mathbf{x}\|_{2}^{2}\bigr{)}}{Z(\mathbf{x})}.\end{split}$ (5)
The above distribution could be factorized coordinatewisely, i.e., $q(\mathbf{x}^{\prime}|\mathbf{x})=\prod_{i=1}^{N}q(\mathbf{x}^{\prime}_{i}|\mathbf{x})$, into a set of categorical distributions:
$q(\mathbf{x}^{\prime}_{i}|\mathbf{x})\propto\exp\biggl{(}{\frac{1}{2}s(\mathbf{x})_{i}(\mathbf{x}_{i}^{\prime}-\mathbf{x}_{i})-\frac{(\mathbf{x}_{i}^{\prime}-\mathbf{x}_{i})^{2}}{2\alpha}\biggr{)}}.$ (6)
When $\mathbf{x}$ is a binary vector, we can obtain the flipping (changing the value of $\mathbf{x}_{i}$ from 00 to $1$, or $1$ to 00) probability:
$q(\mathbf{x}^{\prime}_{i}=1-\mathbf{x}_{i}|\mathbf{x})=\sigma\biggl{(}\frac{1}{2}s(\mathbf{x})_{i}(1-2\mathbf{x}_{i})-\frac{1}{2\alpha}\biggr{)},$ (7)
where $\sigma(\cdot)$ stands for the sigmoid function.
In particular, it can be shown that the discrete Langevin sampler is a first-order approximation to the locally-balanced proposal (Zanella, 2017):
$q(\mathbf{x}^{\prime}|\mathbf{x})\propto\sqrt{p(\mathbf{x}^{\prime})/p(\mathbf{x})}k(\mathbf{x},\mathbf{x}^{\prime}),$ (8)
where $k(\cdot,\cdot)$ is a symmetric function corresponding to $\exp\Bigl{(}-\frac{\|\mathbf{x}^{\prime}-\mathbf{x}\|_{2}^{2}}{2\alpha}\Bigr{)}$ here, and $\log{p(\mathbf{x}^{\prime})}-\log{p(\mathbf{x})}$ (inside the exponential) is approximated by $s(\mathbf{x})^{T}(\mathbf{x}^{\prime}-\mathbf{x})$.
@@ 4.1 Regularized Langevin Dynamics
Although the discrete Langevin sampler can be directly converted into an SA solver by annealing the temperature $\tau$, it nonetheless frequently gets stuck in local optima (see Figure 1 in Section 5.3). As $\tau$ decreases, the gradient term dominates the transition probabilities, driving the flipping probabilities to zero and confining the next sample to a very small neighborhood of the current solution. In order to effectively avoid this undesired behavior, we propose to regularize the expected Hamming distance, i.e., the number of changed coordinates between the sampled and current solutions. Concretely, we start from a locally‐balanced proposal and impose:
$\begin{split}q(\mathbf{x}^{\prime}|\mathbf{x})\propto&\sqrt{p(\mathbf{x}^{\prime})/p(\mathbf{x})}k(\mathbf{x},\mathbf{x}^{\prime}),\\ \text{s.t.}\quad&\mathbb{E}_{q(\mathbf{x}^{\prime}|\mathbf{x})}[\mathrm{Ham}(\mathbf{x}^{\prime},\mathbf{x})]=d,\end{split}$ (9)
where $\mathrm{Ham}(\cdot,\cdot)$ stands for the Hamming distance and $d$ represents the regularized step size. For simplicity, we always treat $d$ as a positive integer in our design. Note that, $k(\mathbf{x},\mathbf{x}^{\prime})$ does not necessarily remain symmetric (and thus the locally-balanced condition may not be satisfied) once the regularization constraint is enforced; instead, it can be chosen to satisfy the constraint at each state.
Since the exact ratio $p(\mathbf{x}^{\prime})/p(\mathbf{x})$ is typically intractable, we follow the discrete Langevin sampler to substitute its first‐order Taylor series expansion:
$\begin{split}q(\mathbf{x}^{\prime}|\mathbf{x})\propto&\exp\biggl{(}{\frac{1}{2}s(\mathbf{x})^{T}(\mathbf{x}^{\prime}-\mathbf{x})\biggr{)}}k(\mathbf{x},\mathbf{x}^{\prime}),\\ \text{s.t.}\quad&\mathbb{E}_{q(\mathbf{x}^{\prime}|\mathbf{x})}[\mathrm{Ham}(\mathbf{x}^{\prime},\mathbf{x})]=d.\end{split}$ (10)
Empirically, this simple regularization technique substantially reduces the tendency to get stuck in local optima. We therefore call our approach Regularized Langevin Dynamics (RLD) and proceed to describe how it integrates into both SA- and NN-based CO solvers.